SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (September 19, 2006)

Attendees: Debbie Bard, Al Eisner, Arik Kreisel, Selina Li,

Operations

  • PTD cables were removed from GLT on Friday as they are no longer needed.

DCT

  • Work is on-going to understand the TSX damage damage problems, but the boards should still be OK for running, with just some occasional X damage (~5 sec/week) showing up. Obviously, it would be better if there were no damage at all!

Data Quality/Monitoring

  • Are the lumi ratio plots in the release yet? We weren't sure but think not. They should still be made retroactively for run 5.

This page is created by Debbie Bard (djbard@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.